# Maximum gap in complex Ginibre matrices

### v1

@@ 3.1. Palm calculus and negative association
for every measurable $F:\mathbf{N}\to\mathbb{R}$ that is bounded and increasing; the latter condition means that $F(\omega_{1})\leq F(\omega_{2})$ if $\omega_{1}\subset\omega_{2}$.
@@
An important property of determinantal point processes is that they are negatively associated (NA) (see [39, Theorem 3.7]). We say that a point process $\nu$ is NA if for each collection of disjoint sets $B_{1},\dots,B_{m}\in\mathcal{B}^{d}$ and each subset $I\subset\{1,\dots,m\}$ we have that
(10) $\mathsf{Cov}(F(\nu(B_{i}),\,i\in I),G(\nu(B_{i}),\,i\in I^{c}))\leq 0,$
for any real, bounded, and increasing functions $F,G$ (see [11]).
@@
Let $F:\mathbf{N}\to\mathbb{R}$ be bounded and increasing. For all Borel sets $A$ with $\sup_{z\in A}|z|<1$ we have
@@
$\int_{A}\mathbb{E}[F(\xi_{n}^{\sqrt{n}z,\varnothing}\cap B_{r_{c,n}}(\sqrt{n}z)^{c})]\mathbf{L}(\mathrm{d}z)$ $=\mathbb{E}\Big{[}\int_{\sqrt{n}A}F(\xi_{n}\cap B_{r_{c,n}}(z)^{c})\mathds{1}\{(\xi_{n}-\delta_{z})(B_{r_{c,n}}(z))=0\}\xi_{n}(\mathrm{d}z)\Big{]}$
$=\int_{\sqrt{n}A}\mathbb{E}[F(\xi_{n}^{z}\cap B_{r_{c,n}}(z)^{c})\mathds{1}\{\xi_{n}^{z!}(B_{r_{c,n}}(z))=0\}]\rho_{n}(z)\,\mathrm{d}z$
$=n\int_{A}\mathbb{E}[F(\xi_{n}^{\sqrt{n}z}\cap B_{r_{c,n}}(\sqrt{n}z)^{c})\mathds{1}\{\xi_{n}^{\sqrt{n}z!}(B_{r_{c,n}}(\sqrt{n}z))=0\}]\rho_{n}(\sqrt{n}z)\,\mathrm{d}z$
$\geq n\int_{A}\mathbb{E}[F(\xi_{n}^{\sqrt{n}z}\cap B_{r_{c,n}}(\sqrt{n}z)^{c})]\mathbb{P}(\xi_{n}^{\sqrt{n}z!}(B_{r_{c,n}}(\sqrt{n}z))=0)\rho_{n}(\sqrt{n}z)\,\mathrm{d}z$
(11) $=\int_{A}\mathbb{E}[F(\xi_{n}^{\sqrt{n}z}\cap B_{r_{c,n}}(\sqrt{n}z)^{c})]\mathbf{L}(\mathrm{d}z),$
@@
where we have used (7) in the first line, (6) in the second line, a change of variables in third line, and (10) together with the fact that $\xi_{n}^{\sqrt{n}z!}$ is a determinantal process (and thus NA) for the inequality in the fourth line. This gives that $\xi_{n}^{\sqrt{n}z,\varnothing}\cap B_{r_{c,n}}(\sqrt{n}z)^{c}\geq\xi_{n}^{\sqrt{n}z}\cap B_{r_{c,n}}(\sqrt{n}z)^{c}$ for $\mathbf{L}$-almost all $z\in B_{1}(0)$.
@@ 3.2. Fast decay of correlation
For all $B\subset B_{1}(0)$ such that $\sup_{z\in B_{1}(0)}|z|<1$, there exists a constant $c(B)>0$ such that
@@
for some constant $C>0$. We note that
@@
The conclusion now follows from the elementary inequality
$x-1-\ln x>0$
for $x\in(0,1)$ and the fact that $|1-z\bar{w}|$ is uniformly lower bounded for $z,w\in B$. ∎
@@
Fix $p,q\in\mathbb{N}$ and set $m=p+q$. For all $B\subset B_{1}(0)$ such that $\sup_{z\in B_{1}(0)}|z|<1$, there exist constants $n_{0}(B),c(B)>0$ such that for all $x_{1},\dots,x_{p+q}$,
@@
Recall that for any determinantal point process with kernel $K(x,y)$, the restriction of the process to some set $S$ is also determinantal, with kernel $K(x,y)\mathds{1}_{S}(x)\mathds{1}_{S}(y)$. Then the conclusion is an immediate consequence of [6, (3.4.5)] and Lemma 3. ∎
@@ 3.3. Infinite Ginibre process
Let $\xi$ be the (stationary) infinite Ginibre process on $\mathbb{C}$ with kernel $K(z,w)=\frac{1}{\pi}e^{-\frac{1}{2}|z|^{2}-\frac{1}{2}|w|^{2}+z\bar{w}}$. Since $K_{n}(z,w)\to K(z,w)$ as $n\to\infty$, we have that $\xi_{n}\to\xi$ in distribution (see, e.g., [16, Theorem 11.1.VII]).
The following result is essentially a well-known theorem of Kostlan [33, Theorem 1.1], but we take the statement from [8, Theorem 4.3.10] to match our choice of scaling. Let $(X_{k})_{k=1}^{\infty}$ be a collection of independent random variables such that $X^{2}_{k}\sim\operatorname{Gamma}(k,1)$, where we use a shape–scale parameterization for the gamma distribution.
@@ 4. Auxiliary statements
Let $s<1$. There exists $C(s)>1$ and $\delta(s)<1$ with the following property. For all $n\in\mathbb{N}$ such that $n\geq C$ and any sequence of measurable sets $A_{n}\subset B_{s\sqrt{n}}$, we have
@@
$\mathbb{P}(\xi_{m}(A_{n})=0)$ $=\frac{1}{m!}\int_{A_{n}^{c}}\dots\int_{A_{n}^{c}}\det(K(z_{i},z_{j}))_{i,j=1}^{m}\,dz_{1}\dots dz_{m}$
$=\frac{1}{m!}\int_{A_{n}^{c}}\dots\int_{A_{n}^{c}}\det(\varphi_{k}(z_{i}))_{i,k=1}^{r}\det(\overline{\varphi_{k}(z_{i})})_{i,k=1}^{m}\,dz_{1}\dots dz_{m}$
@@
[…]
@@
[…]
so that we have $\mathbb{P}(\xi_{m}(A_{n})=0)=\det(M_{m}(A_{n}))$. Since $M_{m}(A_{n})$ in the integral of a positive-definite matrix function, it too is positive-definite. By Fischer’s inequality [31, Theorem 7.8.5],
[…]
@@
where $M_{m}(A_{n})/M_{n}(A_{n})$ is the Schur complement of the block $M_{n}(A_{n})$ of the matrix $M_{m}(A_{n})$. By our assumption that $A_{n}\subset B_{s\sqrt{n}}$, we have $M_{m}(A_{n})\succeq M_{m}(B_{s\sqrt{n}})$. Then, by the monotonicity of the Schur complement [31, 7.7.P41],
@@
[…]
is the probability that $R_{k}>s\sqrt{n}$, where $R_{k}^{2}$ follows a $\operatorname{Gamma}(k,1)$-distribution. For $\lambda>0$, let $Y_{\lambda}$ be Poisson($\lambda$)-distributed and note that $\mathbb{P}(R_{k}^{2}\leq t)=\mathbb{P}(Y_{t}\geq k)$. Hence,
@@
$\prod_{k=n+1}^{\infty}\int_{B_{s\sqrt{n}}^{c}}|\varphi_{k}(w)|^{2}\mathrm{d}w\geq\exp\Big{(}-\sum_{k=n+1}^{\infty}\frac{s_{2}^{k}}{1-s_{2}^{k}}\Big{)}\geq\exp\Big{(}-\frac{s_{2}^{n+1}}{(1-s_{2}^{n})(1-s_{2})}\Big{)}.$
@@
Now we can deduce the asymptotic scaling of the sequence $(r_{n})_{n\in\mathbb{N}}$.
@@
For every Borel set $B$ with $\sup_{z\in B}|z|<1$ we have
@@
Moreover, by the definition of $r_{n}(z)$,
(18) $n\mathbb{P}(\xi_{n}^{z!}(B_{r_{n}}(z))=0)\rho_{n}(z)=c,\quad z\in B_{\sqrt{n}}(0),$
@@
Moreover, from [8, Proposition 7.2.1] we have
@@
$\mathbb{P}(\xi(B_{v_{n}}(0))=0)=\sup_{z\in\sqrt{n}B}\mathbb{P}(\xi(B_{r_{n}(z)}(0))=0)$
@@
To derive the complementary bound, we first estimate the vacuum probability under the reduced Palm measure $\xi_{n}^{z!}$. By [40, Theorem 1] there exists a coupling $(\xi^{\prime},\xi^{\prime\prime})$ of $\xi_{n}^{z!}$ and $\xi_{n}$ such that $\xi^{\prime}\subset\xi^{\prime\prime}$ and $|\xi^{\prime\prime}\setminus\xi^{\prime}|\leq 1$. This gives
@@ 5. Proofs of Main Results
In the next proof, we will need a slight modification of [9, Theorem 3.1], which provides an estimate on the KR distance between two point processes. Let $B\subset B_{1}$. The proof of [9, Theorem 3.1], specialized to our setting, yields
(24) $\begin{split}{d_{\mathrm{KR}}}(\Xi_{n}\cap B,\zeta\cap B)&\leq\sup_{h\in\text{Lip}}\int_{0}^{\infty}\big{|}\mathbb{E}\{\mathcal{L}P_{s}h(\Xi_{n}\cap B)\}\big{|}\mathrm{d}s\\ &=\sup_{h\in\text{Lip}}\int_{0}^{\infty}\left|\int_{B}\mathbb{E}\big{[}D_{z}P_{s}h(\Xi_{n}\cap B)\big{]}-\mathbb{E}\big{[}D_{z}P_{s}h(\Xi_{n}^{z!}\cap B)\big{]}\mathbf{L}(\mathrm{d}z)\right|\mathrm{d}s.\end{split}$
Here $\mathcal{L}$ is the generator of the Glauber dynamics for the Poisson process $\zeta$ (see [9, (2.7)]), and $D_{z}h(\omega)=h(\omega+\delta_{z})-h(\omega)$. We define $T_{n,z}=B_{\log n/\sqrt{n}}(z)$ for $z\in\mathbb{C}$ and set $S_{n,z}=B\setminus T_{n,z}$. We have
(25) $\begin{split}&\Big{|}\mathbb{E}\big{[}D_{z}P_{s}h(\Xi_{n}\cap B)\big{]}-\mathbb{E}\big{[}D_{z}P_{s}h(\Xi_{n}^{x!}\cap B)\big{]}\Big{|}\\ &=\left|\mathbb{E}\left[P_{s}h(\Xi_{n}\cap B+\delta_{z})-P_{s}h(\Xi_{n}^{z!}\cap B+\delta_{z})+P_{s}h(\Xi_{n}\cap B)-P_{s}h(\Xi_{n}^{z!}\cap B)\right]\right|\end{split}$
By [9, (2.9)],
$\mathbb{E}\left[\big{|}P_{s}h(\Xi_{n}\cap B)-P_{s}h(\Xi_{n}\cap S_{n,z})\big{|}\right]$ $\leq e^{-s}\mathbb{E}\big{[}\Xi_{n}(S_{n,z}^{c})\big{]},$
$\mathbb{E}\left[\big{|}P_{s}h(\Xi^{z!}_{n}\cap B)-P_{s}h(\Xi^{z!}_{n}\cap S_{n,z})\big{|}\right]$ $\leq e^{-s}\mathbb{E}\big{[}\Xi^{z!}_{n}(S_{n,z}^{c})\big{]},$
$\mathbb{E}\left[\big{|}P_{s}h(\Xi_{n}\cap B+\delta_{z})-P_{s}h(\Xi_{n}\cap S_{n,z}+\delta_{z})\big{|}\right]$ $\leq e^{-s}\mathbb{E}\big{[}\Xi_{n}(S_{n,z}^{c})\big{]},$
$\mathbb{E}\left[\big{|}P_{s}h(\Xi^{z!}_{n}\cap B+\delta_{z})-P_{s}h(\Xi^{z!}_{n}\cap S_{n,z}+\delta_{z})\big{|}\right]$ $\leq e^{-s}\mathbb{E}\big{[}\Xi^{z!}_{n}(S_{n,z}^{c})\big{]}.$
Then
(26) $\begin{split}&\Big{|}\mathbb{E}\big{[}D_{z}P_{s}h(\Xi_{n}\cap B)\big{]}-\mathbb{E}\big{[}D_{z}P_{s}h(\Xi_{n}^{z!}\cap B)\big{]}\Big{|}\\ &\leq\left|\mathbb{E}\left[P_{s}h(\Xi_{n}\cap S_{n,z})-P_{s}h(\Xi_{n}^{z!}\cap S_{n,z})\right]\right|\\ &+\left|\mathbb{E}\left[P_{s}h(\Xi_{n}\cap S_{n,z}+\delta_{z})-P_{s}h(\Xi_{n}^{z!}\cap S_{n,z}+\delta_{z})\right]\right|+2e^{-s}\left(\mathbb{E}\big{[}\Xi_{n}(S_{n,z}^{c})\big{]}+\mathbb{E}\big{[}\Xi^{z!}_{n}(S_{n,z}^{c})\big{]}\right).\end{split}$
We now specify a coupling of $\Xi_{n}\cap S_{n,z}$ and $\Xi_{n}^{z!}\cap S_{n,z}$ that will allow us to bound the first two terms on the right-hand side of (26). We have by (9) and (11) that $P_{n}^{\sqrt{n}z!}\leq P_{n}$, and $P_{n}^{\sqrt{n}z!}\leq P_{n}^{\sqrt{n}z!,\varnothing}$ on $B_{r_{n}}(\sqrt{n}z)^{c}$ for all $z\in B$. Let $\xi_{n}^{z}\sim P_{n}^{z}$. By Strassen’s theorem (see [35]), there exist point processes $\xi_{n}\sim P_{n}$ and $\hat{\xi}_{n}^{\sqrt{n}z,\varnothing}\sim P_{n}^{\sqrt{n}z,\varnothing}|_{{B_{r_{n}}(\sqrt{n}z)^{c}}}$ such that, almost surely,
$\xi_{n}^{\sqrt{n}z!}\subset\xi_{n},\qquad\xi_{n}^{\sqrt{n}z!}\cap B_{r_{n}}(\sqrt{n}z)^{c}\subset\hat{\xi}_{n}^{\sqrt{n}z!,\varnothing}.$
Since $B_{r_{n}}(w)\subset B_{r_{n}}(\sqrt{n}z)^{c}$ for all $w\in\sqrt{n}S_{n,z}$ by Lemma 8, it holds that $\Xi_{n}[\hat{\xi}_{n}^{\sqrt{n}z,\varnothing}]\cap S_{n,z}$ and $\Xi_{n}[\xi_{n}^{\sqrt{n}z,\varnothing}]\cap S_{n,z}$ agree in distribution, where $\xi_{n}^{\sqrt{n}z,\varnothing}\sim P_{n}^{\sqrt{n}z,\varnothing}$. Therefore, it follows from the definition of $P_{n}^{\sqrt{n}z,\varnothing}$ that $\Xi_{n}[\hat{\xi}_{n}^{z,\varnothing}]\cap S_{n,z}$ is a Palm version of $\Xi_{n}$ at $z$, restricted to $S_{n,z}$. We let $\Xi_{n}^{z}$ denote this version of the process in any expression requiring a coupling below (omitting the restriction to $S_{n,z}$ to lighten the notation).
Returning to (26), we have by our coupling assumption and [9, (2.9)] that
$\left|\mathbb{E}\left[P_{s}h(\Xi_{n}\cap S_{n,z})-P_{s}h(\Xi_{n}^{z!}\cap S_{n,z})\right]\right|$ $\leq e^{-s}\mathbb{E}[(\Xi_{n}\Delta\Xi_{n}^{z!})(S_{n,z})]$
$\left|\mathbb{E}\left[P_{s}h(\Xi_{n}\cap S_{n,z}+\delta_{z})-P_{s}h(\Xi_{n}^{z!}\cap S_{n,z}+\delta_{z})\right]\right|$ $\leq e^{-s}\mathbb{E}[(\Xi_{n}\Delta\Xi_{n}^{z!})(S_{n,z})],$
where we recall that $\Xi_{n}^{z!}:=\Xi^{z}_{n}-\delta_{z}$. Inserting this into (26), and integrating over $s$ in (24), we find that [^3] Note that $z\mapsto\mathbb{E}[(\Xi_{n}\Delta\Xi_{n}^{z!})$ may not be measurable, but our upper bound for it certainly is.

### v2

@@ 3.2. Stochastic domination
for every measurable $F\colon\mathbf{N}\to\mathbb{R}$ that is bounded and increasing; the latter condition means that $F(\omega_{1})\leq F(\omega_{2})$ if $\omega_{1}\subset\omega_{2}$.
@@
An important property of determinantal point processes is that they have negative associations (see [42, Theorem 3.7]). We say that a point process $\nu$ has negative associations if
(15) ${\mathbb{E}[F(\nu)G(\nu)]\leq\mathbb{E}[F(\nu)]\mathbb{E}[G(\nu)],}$
for any real, bounded, and increasing functions $F,G\colon\mathbf{N}\to\mathbb{R}$ that are measurable with respect to complementary subsets (see [37]).
We also require the following lemma concerning stochastic domination.
Lemma 7.
For every $s\in(0,1)$ and $\kappa>1$, there exists $C(s,\kappa)>1$ such that for all $n\geq C$, we have for $L_{n}$-almost all $z\in B_{s}(0)$ that
(16) $P_{n}^{\sqrt{n}z,\varnothing}|_{B_{r_{n}}(\sqrt{n}z)^{c}}\geq P_{n}^{\sqrt{n}z}|_{B_{r_{n}}(\sqrt{n}z)^{c}}.$
Proof.
@@
Let $F\colon\mathbf{N}\to\mathbb{R}$ be bounded and increasing. For all Borel sets $A$ with $\sup_{z\in A}|z|<s$ we have
@@
$\int_{A}\mathbb{E}[F(\xi_{n}^{\sqrt{n}z,\varnothing}\cap B_{r_{n}}(\sqrt{n}z)^{c})]L_{n}(\mathrm{d}z)$ $=\mathbb{E}\left[\int_{\sqrt{n}A}F(\xi_{n}\cap B_{r_{n}}(z)^{c})\mathds{1}\{(\xi_{n}-\delta_{z})(B_{r_{n}}(z))=0\}\xi_{n}(\mathrm{d}z)\right]$
$=\int_{\sqrt{n}A}\mathbb{E}\big[F(\xi_{n}^{z}\cap B_{r_{n}}(z)^{c})\mathds{1}\{\xi_{n}^{z!}(B_{r_{n}}(z))=0\}\big]\rho_{n}(z)\,\mathrm{d}z$
$=n\int_{A}\mathbb{E}\big[F(\xi_{n}^{\sqrt{n}z}\cap B_{r_{n}}(\sqrt{n}z)^{c})\mathds{1}\{\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))=0\}\big]\rho_{n}(\sqrt{n}z)\,\mathrm{d}z$
$\geq n\int_{A}\mathbb{E}\big[F(\xi_{n}^{\sqrt{n}z}\cap B_{r_{n}}(\sqrt{n}z)^{c})\big]\mathbb{P}(\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))=0)\rho_{n}(\sqrt{n}z)\,\mathrm{d}z$
$=\int_{A}\mathbb{E}\big[F(\xi_{n}^{\sqrt{n}z}\cap B_{r_{n}}(\sqrt{n}z)^{c})\big]L_{n}(\mathrm{d}z),$
@@
where we have used (8), Lemma 5, and a change of variables in the first line, (7) in the second line, a change of variables in third line, and (15) together with the fact that $\xi_{n}^{\sqrt{n}z!}$ is a determinantal process (and thus has negative associations) for the inequality in the fourth line. The fifth line follows from (3) and Lemma 5. This gives the claim. ∎
3.3. Coupling of $P_{n}$ and $P_{n}^{z!}$
For a fixed $z\in\mathbb{C}$, we now specify a coupling of $P_{n}$ and $P_{n}^{z!}$. This will help us bound the KR distance of $\Xi_{n}$ and a certain Poisson process in Section 5. By [43, Theorem 1], there exists a coupling $(\xi_{n},\xi_{n}^{z!})$ of $\xi_{n}$ and $\xi_{n}^{z!}$ such that $\xi_{n}^{z!}\subset\xi_{n}$ and $|\xi_{n}^{z!}\setminus\xi_{n}|\leq 1$ almost surely. The hypotheses of [43, Theorem 1] are straightforward to check using the representation (1). We will always use this coupling when discussing the relationship between $\xi_{n}$ and $\xi_{n}^{z!}$ in what follows. Moreover, [43, Theorem 1] also explicitly gives the conditional distribution of $\xi_{n}$ conditional on $\xi_{n}^{z!}$.
Similarly, for any distinct $w,z\in\mathbb{C}$, there exists a coupling of $\xi_{n}^{z!}$ and $\xi_{n}^{z!,w!}:=(\xi_{n}^{z!})^{w!}$. The hypotheses of [43, Theorem 1] may be checked in this case using (10).
Additionally, under the conditions of Lemma 7, we specify a coupling of $P_{n}^{\sqrt{n}z!}$ and $P_{n}^{\sqrt{n}z!,\varnothing}$ on $B_{r_{n}}(\sqrt{n}z)^{c}$. We use in coupling in Section 5 to construct a coupling of $\Xi_{n}$ and $\Xi_{n}^{z!}$ outside of a ball around $z$. We do not use [43, Theorem 1] for this purpose, since its hypotheses are harder to check in this case, and we instead settle for a weaker statement. Fix $z\in B_{1}(0)$ and $\xi_{n}^{\sqrt{n}z!}\sim P_{n}^{\sqrt{n}z!}$, and observe (16) gives that $P_{n}^{\sqrt{n}z!}\leq P_{n}^{\sqrt{n}z!,\varnothing}$ on $B_{r_{n}}(\sqrt{n}z)^{c}$ (for sufficiently large $n$). By Strassen’s theorem (see [38]), there exists a measure $\rho\sim P_{n}^{\sqrt{n}z!}\otimes P_{n}^{\sqrt{n}z!,\varnothing}$ supported on $\{\omega_{1}\otimes\omega_{2}\in\mathbf{N}\times\mathbf{N}:\omega_{1}\subset\omega_{2}\}$. By [34, Theorem 3.4(i)], $\rho$ can be disintegrated and, hence, there exists a point process $\xi_{n}^{z!,\varnothing}\sim P_{n}^{\sqrt{n}z!,\varnothing}$ such that, almost surely,
(17) $\xi_{n}^{\sqrt{n}z!}|_{B_{r_{n}}(\sqrt{n}z)^{c}}\subset\xi_{n}^{\sqrt{n}z!,\varnothing}|_{B_{r_{n}}(\sqrt{n}z)^{c}}.$
Combining the above couplings gives a joint coupling of $\xi_{n}|_{B_{r_{n}}(\sqrt{n}z)^{c}}$, $\xi_{n}^{\sqrt{n}z!}|_{B_{r_{n}}(\sqrt{n}z)^{c}}$, and $\xi_{n}^{\sqrt{n}z!,\varnothing}|_{B_{r_{n}}(\sqrt{n}z)^{c}}$ such that $\xi_{n}^{\sqrt{n}z!}|_{B_{r_{n}}(\sqrt{n}z)^{c}}\subset\xi_{n}|_{B_{r_{n}}(\sqrt{n}z)^{c}}$ and (17) hold.
@@ 3.4. Fast decay of correlation
For all $B\subset B_{1}(0)$ such that $\sup_{z\in B}|z|<1$, there exists a constant $c(B)>0$ such that
@@
for some constant $C>1$. We note that
@@
The conclusion now follows from the elementary fact that there exists $c(B)>0$ such that
$x-1-\ln x>c$
for $x$ such that $0\leq x\leq\sup_{z\in B}|z|$, and the fact that $|1-z\bar{w}|$ is uniformly lower bounded for $z,w\in B$. ∎
The following lemma concerns a mixing property of $\xi_{n}$.
@@
Let $\rho^{(m)}$ be the $m$-th order correlation function of $\xi_{n}$. Fix $p,q\in\mathbb{N}$ and set $m=p+q$. For all $B\subset B_{1}(0)$ such that $\sup_{z\in B_{1}(0)}|z|<1$, there exist constants $n_{0}(B),c(B)>0$ such that for all $x_{1},\dots,x_{p+q}$,
@@
Recall that for any determinantal point process with kernel $K(x,y)$, the restriction of the process to some set $S$ is also determinantal, with kernel $K(x,y)\mathds{1}_{S}(x)\mathds{1}_{S}(y)$. Then the conclusion is an immediate consequence of [7, (3.4.5)] and Lemma 8. ∎
@@ 3.5. Infinite Ginibre process
Let $\xi$ be the (stationary) infinite Ginibre process on $\mathbb{C}$ with kernel
$K(z,w)=\frac{1}{\pi}e^{-\frac{1}{2}|z|^{2}-\frac{1}{2}|w|^{2}+z\bar{w}}.$
Since $K_{n}(z,w)\to K(z,w)$ as $n\to\infty$, we have that $\xi_{n}\to\xi$ in distribution (see, e.g., [17, Theorem 11.1.VII]).
The following result is essentially a well-known theorem of Kostlan [35, Theorem 1.1], but we take the statement from [9, Theorem 4.3.10] to match our choice of scaling. Let $(X_{k})_{k=1}^{\infty}$ be a collection of independent random variables such that $X^{2}_{k}\sim\operatorname{Gamma}(k,1)$, where we use a shape–scale parameterization for the gamma distribution.
@@ 4. Auxiliary statements
For every $s\in(0,1)$, there exists $C(s)>1$ and $\delta(s)<1$ with the following property. For all $n\in\mathbb{N}$ such that $n\geq C$ and any sequence of measurable sets $A_{n}\subset B_{s\sqrt{n}}$, we have
@@
$\mathbb{P}(\xi_{m}(A_{n})=0)$ $=\frac{1}{m!}\int_{A_{n}^{c}}\dots\int_{A_{n}^{c}}\det(K_{m}(z_{i},z_{j}))_{i,j=1}^{m}\,dz_{1}\dots dz_{m}$
$=\frac{1}{m!}\int_{A_{n}^{c}}\dots\int_{A_{n}^{c}}\det(\varphi_{k}(z_{i}))_{i,k=1}^{m}\det(\overline{\varphi_{k}(z_{i})})_{i,k=1}^{m}\,dz_{1}\dots dz_{m}$
@@
$=\sum_{\sigma\in S_{n}}\operatorname{sgn}(\sigma)\prod_{i=1}^{m}\int_{A_{n}^{c}}\varphi_{i}(z)\overline{\varphi_{\sigma(i)}(z)}\,dz$
$=\det\left(\int_{A_{n}^{c}}\varphi_{i}(w)\overline{\varphi_{j}(w)}dw\right)_{1\leq i,j\leq m},$
@@
$M_{m}(U):=\Big(\int_{U^{c}}\varphi_{i}(w)\overline{\varphi_{j}(w)}dw\Big)_{1\leq i,j\leq m},$
so that we have $\mathbb{P}(\xi_{m}(A_{n})=0)=\det(M_{m}(A_{n}))$. For every $U\subset\mathbb{C}$, $M_{m}(U)$ is the integral of a positive-definite matrix function, so it too is positive-definite. In particular $M_{m}(A_{n})$ is positive-definite, so by Fischer’s inequality [32, Theorem 7.8.5],
$\det(M_{m}(A_{n}))\leq\det(M_{m-1,n})\left(\int_{A_{n}^{c}}\varphi_{m}(w)\overline{\varphi_{j}(w)}dw\right)\leq\det(M_{m-1,n}),$
@@
where $M_{m}(A_{n})/M_{n}(A_{n})$ is the Schur complement of the block $M_{n}(A_{n})$ of the matrix $M_{m}(A_{n})$. By our assumption that $A_{n}\subset B_{s\sqrt{n}}$, we have $M_{m}(A_{n})\succeq M_{m}(B_{s\sqrt{n}})$, since
$M_{m}(A_{n})-M_{m}(B_{s\sqrt{n}})=M_{m}\big(A_{n}\cup B_{s\sqrt{n}}^{c}\big)\succeq 0,$
where $A\succeq B$ means that ABA-B is positive-definite for $m\times m$-matrices $A,B$. Then, by the monotonicity of the Schur complement [32, 7.7.P41],
@@
$\int_{B_{s\sqrt{n}}^{c}}|\varphi_{k}(w)|^{2}\mathrm{d}w=\frac{1}{\pi(k-1)!}\int_{B_{s\sqrt{n}}^{c}}|w|^{2(k-1)}e^{-|w|^{2}}\,\mathrm{d}w=\frac{2}{(k-1)!}\int_{s\sqrt{n}}^{\infty}r^{2k-1}e^{-r^{2}}\,dr$
is the probability that $R_{k}>s\sqrt{n}$, where $R_{k}^{2}$ follows a $\operatorname{Gamma}(k,1)$-distribution. For all $\lambda>0$, let $Y_{\lambda}$ be Poisson($\lambda$)-distributed and note that $\mathbb{P}(R_{k}^{2}\leq t)=\mathbb{P}(Y_{t}\geq k)$. Hence,
@@
$\prod_{k=n+1}^{\infty}\int_{B_{s\sqrt{n}}^{c}}|\varphi_{k}(w)|^{2}\mathrm{d}w\geq\exp\left(-\sum_{k=n+1}^{\infty}\frac{s_{2}^{k}}{1-s_{2}^{k}}\right)\geq\exp\left(-\frac{s_{2}^{n+1}}{(1-s_{2}^{n})(1-s_{2})}\right).$
@@
Now we can deduce the asymptotic scaling of the sequence $(r_{n})_{n\in\mathbb{N}}$. We recall that $r_{n}$ depends on a parameter $\kappa>0$ that we suppress in the notation.
@@
For every Borel set $B$ with $\sup_{z\in B}|z|<1$, and all $\kappa>1$, we have
@@
Moreover, by Lemma 5, there exists $C(B,\kappa)>1$ such that for $n\geq C$,
(24) $n\mathbb{P}(\xi_{n}^{z!}(B_{r_{n}}(z))=0)\rho_{n}(z)=\kappa,\quad z\in B_{\sqrt{n}}(0),$
@@
Moreover, from [9, Proposition 7.2.1] we have
@@
$\mathbb{P}(\xi(B_{v_{n}}(0))=0)=\sup_{z\in\sqrt{n}B}\mathbb{P}(\xi(B_{r_{n}(z)}(0))=0),$
@@
To derive the complementary bound, we first estimate the vacuum probability under the reduced Palm measure $\xi_{n}^{z!}$. Using the coupling of $\xi_{n}$ and $\xi_{n}^{z!}$ introduced in Section 3.3, we have
@@ 5. Proofs of Main Results
In the next proof, we will need a slight modification of [10, Theorem 3.1], which provides an estimate on the KR distance between two point processes that is based on Stein’s method. Since we are unable to verify that $\mathbb{E}[(\Xi_{n}[\xi_{n}]\Delta\Xi_{n}[\xi_{n}^{z!,\varnothing}])(B)]$ is measurable as a function of $z$ for an arbitrary $B\subset B_{1}(0)$, we cannot apply [10, Theorem 3.1] directly to obtain the desired result. Instead, we adapt the arguments given there to our setting.
Lemma 14.
Fix $\kappa>1$, and let $\zeta_{\kappa}$ be a Poisson process with intensity measure $L_{n}$. Fix $t\in(0,1)$. There exists a constant $C(t,\kappa)>1$ such that for all $B\subset B_{t}(0)$ and $n\geq C$, we have